OpenStudy (anonymous):

Evaluate -3x^3 - 4x for x = -1. 1 7 -1

OpenStudy (alexandervonhumboldt2):

-3*(-1)^3+4=-3*-1+4=3+4=7

OpenStudy (alexandervonhumboldt2):

7

OpenStudy (anonymous):

thank you so much!! =))))

OpenStudy (alexandervonhumboldt2):

can you plz medal?